A group of students got the following scores in a test: 6,9,13,16,18,21. Consider sample size of size 3 that can be drawn from this population.




A. List all the possible samples and the corresponding mean.




B. Construct the samplong distribution of the sample means.




C. Compute for the mean and standard deviation of the sampling distribution of the sample means.

1
Expert's answer
2022-06-01T08:10:36-0400

We have population values 6,9,13,16,18,21, population size N=6 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{6+9+13+16+18+21}{6}=\dfrac{83}{6}"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{216}(2209+841+25""+169+625+1849)=\dfrac{953}{36}""\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{953}{36}}\approx5.1451"


A. Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_3=20."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 6,9,13 & 28/3 \\
 \hdashline
 2 & 6,9,16 & 31/3 \\
 \hdashline
 3 & 6,9,18 & 33/3 \\
 \hdashline
 4 & 6,9,21 & 36/3 \\
 \hdashline
 5 & 6,13,16 & 35/3 \\
 \hdashline
 6 & 6,13,18 & 37/3 \\
 \hdashline
 7 & 6,13, 21 & 40/3 \\
 \hdashline
 8 & 6,16,18 & 40/3 \\
 \hdashline
 9 & 6,16,21 & 43/3 \\
 \hdashline
 10 & 6, 18,21 & 45/3 \\
 \hdashline
 11 & 9,13,16 & 38/3 \\
 \hdashline
 12 & 9, 13,18 & 40/3 \\
 \hdashline
 13 & 9, 13, 21 & 43/3 \\
 \hdashline
 14 & 9,16,18 & 43/3 \\
 \hdashline
 15 & 9,16,21 & 46/3 \\
 \hdashline
 16 & 9, 18,21 & 48/3 \\
 \hdashline
 17 & 13, 16,18 & 47/3 \\
 \hdashline
 18 & 13, 16, 21 & 50/3 \\
 \hdashline
 19 & 13, 18,21 & 52/3 \\
 \hdashline
 20 & 16, 18,21 & 55/3 \\
 \hdashline
\end{array}"



B.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 28/3 & 1/20 & 28/60 & 784/180 \\
 \hdashline
 31/3 & 1/20 & 31/60 & 961/180 \\
 \hdashline
 33/3 & 1/20 & 33/60 & 1089/180 \\
 \hdashline
 35/3 & 1/20 & 35/60 & 1225/180 \\
 \hdashline
 36/3 & 1/20 & 36/60 & 1296/180 \\
 \hdashline
 37/3 & 1/20 & 37/60 & 1369/180 \\
 \hdashline
 38/3 & 1/20 & 38/60 & 1444/180 \\
 \hdashline
 40/3 & 3/20 & 120/60 & 4800/180 \\
 \hdashline
 43/3 & 3/20 & 129/60 & 5547/180 \\
 \hdashline
 45/3 & 1/20 & 45/60 & 2025/180 \\
 \hdashline
 46/3 & 1/20 & 46/60 & 2116/180 \\
 \hdashline
 47/3 & 1/20 & 47/60 & 2209/180 \\
 \hdashline
 48/3 & 1/20 & 48/60 & 2304/180 \\
 \hdashline
 50/3 & 1/20 & 50/60 & 2500/180 \\
 \hdashline
 52/3 & 1/20 & 52/60 & 2704/180 \\
 \hdashline
 55/3 & 1/20 & 55/60 & 3025/180 \\
 \hdashline
\end{array}"



Mean of sampling distribution 



"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{830}{60}=\dfrac{83}{6}=\mu"



The variance of sampling distribution 



"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{35398}{180}-(\dfrac{83}{6})^2=\dfrac{953}{180}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})""\sigma_{\bar{X}}=\sqrt{\dfrac{953}{180}}\approx2.3010"